OpenStudy (anonymous):

Find the equation of a line parallel to x-2y=6 and has the same x- intercept as 4x + 3y=-12.

OpenStudy (anonymous):

1) x-2y=6 -2y=6-x y=1/2 x -3 so, the slope of your line should be = 1/2 (because parallel)

OpenStudy (anonymous):

2) x-intercept of the second equation means y=0, so 4x+3*0=-12 x=-12/4=-3 so, the coordinate of the intercept: (-3,0)

OpenStudy (anonymous):

3) now we build the equation: y-0=1/2 (x-(-3)) y=1/2 (x+3) y=1/2 x +3/2

OpenStudy (anonymous):

thank you so much inik :)